OpenStudy (anonymous):

Find the equation of the tangent line to the curve (a lemniscate) 2(x^2+y^2)2=25(x^2−y^2) at the point (−3,1). The equation of this tangent line can be written in the form y=mx+b where m is:

OpenStudy (anonymous):

2) whre b is

OpenStudy (anonymous):

Find y', evalute it at your point to get the slope of the tangent line at the point of interest. Write the line in point-slope form, convert to slope-intercept form.

OpenStudy (anonymous):

Do i need to caculate the equation then find y'?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

do i need to simplif first

OpenStudy (anonymous):

Differentiate implicitly. Do your simplification when you are trying to isolate y'. It's not that messy

OpenStudy (anonymous):

is the first part: 2(x^2-y^2)2? If so, that's the same as saying 4(x^2 - y^2) @Dodo1 Unless you made a typing mistake..

OpenStudy (anonymous):

I assumed that it's squared because lemniscate

OpenStudy (anonymous):

\[f(x)=2(x^2+y^2)^2=25(x^2-y^2)\]

OpenStudy (anonymous):

sorry I typed it wrong

OpenStudy (anonymous):

So Do i need to similfy than Y'?

OpenStudy (anonymous):

or move y first?

OpenStudy (anonymous):

Just differentiate implicitly like I said

OpenStudy (anonymous):

mmmm how do you do that

OpenStudy (anonymous):

Is it product rule that I am going to use

OpenStudy (anonymous):

?